OpenStudy (anonymous):

find dy/dx at x=0 y=u-(2/u) u= (3x+1)^3

OpenStudy (amistre64):

dy/du = 1+(2/u^2) du/dx = 3*3(3x+1)^2

OpenStudy (amistre64):

2 9(3x+1)^2 (1+ -------- ) * ( ---------) (3x+1)^6 1

OpenStudy (amistre64):

18(3x+1)^2 9(3x+1)^2 + ------------ (3x+1)^6 9(3x+1)^6 + 18 --------------- ?? (3x+1)^4

OpenStudy (anonymous):

how did you get dy/du =u-(2/u)?

OpenStudy (amistre64):

y = u -(2/u) dy = 1 - (-2/u^2) du dy/du = 1 + (2/u^2)

OpenStudy (amistre64):

at x=0; id say that i did the calculation right and go with: 27

OpenStudy (anonymous):

ohh got it! thanks!

OpenStudy (amistre64):

you could simply plug the value of u into the equation for f(u)... but the chain rule makes life easier